OpenStudy (anonymous):

Jennifer can clean her room 2 hours, while John can clean the same room in 5 hours. How many hours would it take for them to clean the room together?

OpenStudy (anonymous):

Wouldnt it be just like 1 hour

OpenStudy (anonymous):

more than that but less than 2hr

OpenStudy (anonymous):

my answer choices are wierd. its 7, 10, 7 over 10, or 10 over 7. I dont understand

OpenStudy (anonymous):

10/7

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{1}{5}=\frac{1}{x} \]

OpenStudy (anonymous):

the sum of their rates is 1/2 + 1/5 and so 1 / (1/2 + 1/5) = 10/7

OpenStudy (anonymous):

but where did 10 come from

OpenStudy (anonymous):

well thanks for the help

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{1}{5}=\frac{1}{x} \]\[\frac{7}{10}=\frac{1}{x} \]7x = 10\[x=\frac{10}{7} \]

OpenStudy (anonymous):

1/2 and 1/5 can interpreted as 1 job per 2 hours and 1 job per 5 hours respectively.